OpenStudy (anonymous):

Function notation -x+5y=5

OpenStudy (dtan5457):

First you have to isolate y. Function notation is just f(x)= Simply replacing y with f(x)

OpenStudy (anonymous):

I'm trying it don't wanna work for me lol

OpenStudy (anonymous):

It's due Friday at 12:30pm

OpenStudy (hockeychick23):

-x+5y=5 Ok so first lets add x to both sides -x+x+5y=5+x 5y=x+5 Divide both sides by 5 5y/5=x/5+5 y= 1/5x+1